OpenStudy (anonymous):

Which property could be used to find the missing number? (–7.3 + 5.2) + 11 = –7.3 + ( + 11) A. commutative property B. associative property C. distributive property

OpenStudy (puppylover52):

can u help me

OpenStudy (lalipapa123):

comming!

OpenStudy (lalipapa123):

what do you think the asnwer was?

OpenStudy (anonymous):

A or C not sure @iGreen Need lot's of help can you come

OpenStudy (igreen):

Commutative Property: \(\sf a + b = b + a\) Associative Property: \(\sf (a + b) + c = a + (b + c)\) Distributive Property: \(\sf a(b + c) = ab + ac\)

OpenStudy (igreen):

So which one do you think?

OpenStudy (anonymous):

B

OpenStudy (igreen):

Correct!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YES 100% mastered thank you so much @iGreen

OpenStudy (anonymous):

thanks @puppylover52